# Sequential Convex Programming for Multimode Spacecraft Trajectory Optimization

### v1

@@ 2.1 Dynamics and propulsion models
The SEP propulsion model assumes a control vector $\boldsymbol{T}_{i}$ with $\|\boldsymbol{T}_{i}\|\leq 1$ which defines the thrust direction and throttle factor for each thrust mode $i$. This can change over time depending on the thrust transcription strategy; here we assume it is held constant throughout each segment of the trajectory as in a zero-order-hold (ZOH) transcription. The thrust acceleration and mass flow rate for a set of SEP modes can then be expressed as,
@@
where $r_{\oplus}$ is the reference distance from the sun (typically 1 AU), $C$ is the solar flux at $r_{\oplus}$, $A$ is the sail area, $m$ is the spacecraft mass, $r$ is the distance of the spacecraft from the Sun, $\hat{\boldsymbol{u}}_{n}$ is the unit normal vector of the sail, $\hat{\boldsymbol{u}}_{r}$ is the unit vector from the spacecraft to the Sun, $\mu$ is the solar sail specular reflection coefficient, and $\nu$ is the solar sail diffuse reflection coefficient. In this work, the sail normal $\hat{\boldsymbol{u}}_{n}$ is parameterized for output in terms of the cone angle $\alpha$ and clock angle $\beta$, which define the orientation of the sail normal with respect to the sunlight direction $\hat{\boldsymbol{u}}_{r}$.
@@ 2.2 Sequential convex programming
Applying the general approach of SCP (?), the trajectory is split into many parts as in a direct method, each referred to as a segment with index n=1,2,,Nn=1,2,...,N. The number of segments $N$ is calculated based on an intended timespan between potential impulses. With this discretization, the linearized dynamic constraints are constructed around a reference trajectory, which requires an appropriate initial guess. For this work, a ballistic reference trajectory with zero controls is used. Then, each segment is assigned a single control input for each propulsion mode, which are held constant throughout the segment as in a ZOH transcription. For SEP propulsion, this is the thrust direction and throttle factor $\boldsymbol{T}_{i}$ for each mode $i$. For solar sailing, it is the sail acceleration vector $\boldsymbol{a}_{\text{SAIL}}$.
@@
$\boldsymbol{d}_{n}$ $=\boldsymbol{\bar{x}}_{n}-\boldsymbol{A}_{n}\boldsymbol{\bar{x}_{n}}-\sum_{i=1}^{N}\boldsymbol{B}_{n,i}\boldsymbol{T}_{n,i}-\boldsymbol{C}_{n}\boldsymbol{a}_{\text{SAIL},n}.$ (9)
@@
Sparse AD (?) is employed to improve the computational efficiency of computing the entire dynamics constraint matrix, improve flexibility and to automatically detect the non-zero entries. Instead of computing each $\boldsymbol{A}$, $\boldsymbol{B}_{i}$ and $\boldsymbol{C}$ matrix for each segment, the entire dynamics constraint matrix can be computed efficiently in the minimal number of calls through matrix coloring and subsequent rearrangement. This matrix rearrangement process allows for the efficient use of forward-mode AD which can compute columns in a single pass. This avoids the need to manually compute and design a discretization scheme for the dynamical constraint, enabling a significant improvement in flexibility and ease of implementation of various dynamical models. The resulting constraint matrix and coloring pattern is illustrated in Fig. 1 for a problem with $N=5$ segments, $1$ SEP mode and $1$ solar sail. The sparsity pattern is clearly visible, and the coloring pattern shows that only $14$ calls of forward-mode AD (one for each color) are required to compute the entire constraint matrix, rather than the 63 calls (one for each column) without coloring.

### v2

@@ 2.1 Dynamics and propulsion models
The SEP propulsion model assumes a control vector $\boldsymbol{T}_{i}$ with $\|\boldsymbol{T}_{i}\|\leq 1$ which defines the thrust direction and throttle factor for each thrust mode ii. This can change over time depending on the thrust transcription strategy; here we assume it is held constant throughout each segment of the trajectory as in a zero-order-hold (ZOH) transcription. The thrust acceleration and mass flow rate for a set of SEP modes can then be expressed as,
@@
where $r_{\oplus}$ is the reference distance from the sun (typically 1 AU), $C$ is the solar flux at $r_{\oplus}$, AA is the sail area, $m$ is the spacecraft mass, $r$ is the distance of the spacecraft from the Sun, $\hat{\boldsymbol{u}}_{n}$ is the unit normal vector of the sail, $\hat{\boldsymbol{u}}_{r}$ is the unit vector from the spacecraft to the Sun, $\mu$ is the solar sail specular reflection coefficient, and $\nu$ is the solar sail diffuse reflection coefficient. In this work, the sail normal $\hat{\boldsymbol{u}}_{n}$ is parameterized for output in terms of the cone angle $\alpha$ and clock angle $\beta$, which define the orientation of the sail normal with respect to the sunlight direction $\hat{\boldsymbol{u}}_{r}$.
@@ 2.2 Sequential convex programming
Applying the general approach of SCP (?), the trajectory is split into many parts as in a direct method, each referred to as a segment with index n=1,2,,Nn=1,2,...,N. The number of segments NN is calculated based on an intended timespan between potential impulses. With this discretization, the linearized dynamic constraints are constructed around a reference trajectory, which requires an appropriate initial guess. For this work, a ballistic reference trajectory with zero controls is used. Then, each segment is assigned a single control input for each propulsion mode, which are held constant throughout the segment as in a ZOH transcription. For SEP propulsion, this is the thrust direction and throttle factor $\boldsymbol{T}_{i}$ for each mode ii. For solar sailing, it is the sail acceleration vector $\boldsymbol{a}_{\text{SAIL}}$.
@@
$\boldsymbol{d}_{n}$ $=\boldsymbol{\bar{x}}_{n+1}-\boldsymbol{A}_{n}\boldsymbol{\bar{x}_{n}}-\sum_{i=1}^{N}\boldsymbol{B}_{n,i}\boldsymbol{T}_{n,i}-\boldsymbol{C}_{n}\boldsymbol{a}_{\text{SAIL},n}.$ (9)
@@
Sparse AD (?) is employed to improve the computational efficiency of computing the entire dynamics constraint matrix, improve flexibility and to automatically detect the non-zero entries. Instead of computing each $\boldsymbol{A}$, $\boldsymbol{B}_{i}$ and $\boldsymbol{C}$ matrix for each segment, the entire dynamics constraint matrix can be computed efficiently in the minimal number of calls through matrix coloring and subsequent rearrangement. This matrix rearrangement process allows for the efficient use of forward-mode AD which can compute columns in a single pass. This avoids the need to manually compute and design a discretization scheme for the dynamical constraint, enabling a significant improvement in flexibility and ease of implementation of various dynamical models. The resulting constraint matrix and coloring pattern is illustrated in Fig. 1 for a problem with $N=5$ segments, 11 SEP mode and 11 solar sail. The sparsity pattern is clearly visible, and the coloring pattern shows that only $14$ calls of forward-mode AD (one for each color) are required to compute the entire constraint matrix, rather than the 63 calls (one for each column) without coloring.